If the angle of a pentagon could be x, 2x, 3x, 4x, and 5x.what would be the value of x
Answer by stanbon(75887)

If the angle of a pentagon could be x, 2x, 3x, 4x, and 5x.what would be the value of x
-----
Sum of the angles = (5-2)180 = 3*180 = 540 degrees
----
Equation:
x + 2x + 3x + 4x + 5x = 540
15x = 540
x = 36
=============
Cheers,
Stan H.
==========